# DIRECT: Deep Active Learning under Imbalance and Label Noise

### v2

@@ Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods. In this work, we conduct the first study of active learning under both class imbalance and label noise. We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise. We present extensive experiments on imbalanced datasets with and without label noise. Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
Machine Learning, ICML
@@ 1 Introduction
Large-scale deep learning models are playing increasingly important roles across many industries. Human feedback and annotations have played a significant role in developing such systems through. Progressively over time, we believe the role of humans in a machine learning pipeline would shift to annotating rare yet important cases. However, under data imbalance, the typical strategy of randomly choosing examples for annotation becomes especially inefficient. This is because the majority of the labeling budget would be spent on common and well-learned classes, resulting in insufficient rare class examples for training an effective model. For many annotation jobs, this challenge of data imbalance is further compounded by label noises induced from annotators. In this paper, we mitigate these issues by proposing a deep active learning strategy. Our algorithm DIRECT sequentially and adaptively chooses informative and more class-balanced examples for annotation while being robust to noisy annotations. To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
@@
Inspired by previous work by Zhang et al. (2022), we propose a novel reduction of the imbalanced classification problem into one-dimensional active learning problems. For each class, our reduction sorts unlabeled examples into an list ordered by one-vs-rest margin scores. The goal of DIRECT is to find the optimal separation threshold which best separates the examples in any given class from the rests. With an equivalence to finding the best threshold classifier, we were able to employ classic active learning literature for obtaining the separation threshold. Once the optimal separation threshold is identified, DIRECT annotates a set of more class-balanced examples near this threshold. As examples around the optimal separation threshold tend to have high uncertainty, DIRECT annotates examples that are both class-balanced and informative.
With the reduction to classic one-dimensional active learning literature, DIRECT also inherits the advantages from classic active learning algorithms such as robustness against label noise and batch labeling. This is in contrast to the state-of-art imbalanced active learning algorithm GALAXY proposed by Zhang et al. (2022), which requires sequential labeling and is much less robust against label noise. As demonstrated in Figure 1, under both noisy and noiseless label settings, DIRECT improves significantly over GALAXY’s label efficiency – DIRECT requires less annotations to achieve the same accuracy. It achieves state-of-art label efficiency gain on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench framework (Zhang et al., 2023a). Lastly, DIRECT also allows for parallel annotations with various batch sizes.
@@
Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings. The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy. Both (a) and (b) are using supervised training of ResNet-18. In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2023a). $B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators for the annotation job. $B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY. Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world decisions.
@@
  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification for deep neural networks.
@@
  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32). DIRECT consistently performs the best compared against existing baseline algorithms. Quantitatively, DIRECT can save more than $60\%$ annotation cost compared to GALAXY, and more than $80\%$ annotation cost compared to random sampling.
@@ 2 Related Work
Class-Balanced Deep Active Learning Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009). Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a). More sophisticated methods have also been proposed to optimize the chosen unlabeled examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as under label noises (Khosla et al., 2022) or under class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2023a).
@@
Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications. Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2023b). To this end, Kothawade et al. (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances. Our work is motivated by Zhang et al. (2022), where they propose GALAXY, an algorithm that construct one-dimensional linear graphs and apply graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain. While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation. In addition, bisection procedures are generally not robust against label noises, a prevalent challenge in real world annotation tasks. Our algorithm DIRECT mitigates all of the above shortcomings of GALAXY while outperforming it even with synchronous labeling and no label noise, beating GALAXY in its own game.
Lastly, we distinguish our work from Zhang et al. (2023b), where the paper studies the algorithm selection problem. Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature. Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021). In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon. This is fundamentally different from our work, where our annotator may provide noisy labels. Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022). The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on those instances. In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today. Recently, Khosla et al. (2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise. Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
To the best of our knowledge, no existing deep active learning work has studied the scenario where both imbalance and label noise present. Yet, this setting is the most prevalent in real-world annotation applications.
@@
Agnostic Active Learning A critical part of our solution is the reduction to a one dimensional agnostic active learning problem, e.g. Section 4.1. There is extensive work on agnostic learning in general, and we refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion. All of these works, beginning with the seminal works (Balcan et al., 2006; Dasgupta et al., 2007) follow, a familiar paradigm of disagreement based learning. This involves maintainin a version space of promising hypothesis at each time, and constructing the disagreement region as a set of unlabeled examples for which there exists two hypotheses that disagree on their predictions. The example to label is chosen by sampling from a informative sampling distribution computed over the disagreement region. Several approaches have been proposed for computing the sampling distribution, e.g. Jain & Jamieson (2019); Katz-Samuels et al. (2020, 2021); Huang et al. (2015). As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al. (2021). We remark that the problem of class imbalance has been largely ignored by the classical work on traditional agnostic learning, and this paper hopes to close this gap.
@@ 3.1 Notations
We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation. Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown. Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted $K$ classes. Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ of each class. We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k^{\prime}\in[K]}N_{k^{\prime}}}$.
@@
3.2 Review of Imbalanced Active Learning: GALAXY
We first consider a binary case of imbalanced classification, where we assume $N_{1}<N_{2}$ without loss of generality. After a total of $T$ rounds and sampling $TB$ examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$. On the other hand, as shown previously in Zhang et al. (2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low. The reasoning is simply demonstrated in Figure 1(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples almost all from the majority class. It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
@@ 3.2 Review of Imbalanced Active Learning: GALAXY
Figure 2: Figures demonstrating GALAXY. Ordered lists of examples are ranked by the predictive sigmoid score $\widehat{p}$. The ground truth label of each example is represented by its border – solid blue lines for class $1$ and dotted red lines for class $2$. Annotated examples are shaded.
Therefore, the key objective in imbalanced active learning is to label examples that are both uncertain and class balanced. In both GALAXY (Zhang et al., 2022) and this work, we decompose the objective into the following two-phased procedure: 1. Identify the optimal separation threshold $j^{\star}$ that best separates the minority class from the majority class, as shown in Figure 1(a). 2. Annotate equal number of examples next to $j^{\star}$ from both sides.
GALAXY draws inspiration from graph-based active learning in identifying $p^{\star}$. In particular, it follows a modified bisection procedure that sequentially annotates examples to find all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 1(b)). As the optimal separation threshold must be located at one of these cuts, GALAXY annotates examples close to all the found cuts. However, the algorithm suffers from three weaknesses. First, as shown in Figure 1(b), by annotating around all cuts, the algorithm may waste a significant portion of the annotation budget around outliers, which could lead to annotating large number of majority class examples. Second, the above problem can be magnified when annotators provide noisy labels. Third, a bisection procedure only allows for sequential labeling, which prevents multiple annotators labeling in parallel. In this paper, we take a DIRECT approach by identifying only the optimal separation threshold and address all of the shortcomings above.
4 A Robust Algorithm with Batch Labeling
@@ 4.1 An 1-D Reduction to Agnostic Active Learning
We start by considering the imbalanced binary classification setting mentioned in Section 3.2. When given a neural network model, we let $\widehat{p}:X\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores. Here, a higher sigmoid score represents a higher confidence of the example being in class $2$. We sort examples by their sigmoid predictive score similar to Section 3.2. Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ be a sorted permutation of $\widehat{p}(x_{1}),...,\widehat{p}(x_{N})$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels. We are now ready to define the optimal separation threshold as described in Section 3.2.
@@
Given a sorting of $X$ the points according the sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
@@
$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=2:i>j\}\rvert\right).$ (1)
@@
In other words, on either side of $j^{\ast}$, it has the largest discrepancy in the number of examples between the two classes. This captures the intuition of Figure 1(b) — our goal is to find a threshold that best separates one class from the other. We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ is defined as
$h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}.$
@@
and $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times. The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1). Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
@@
Multi-Class Classification. To generalize the above problem formulation to multi-class classification, we follow a similar strategy to Zhang et al. (2022). Specifically, for each class $k$, one can view the problem of class-$k$ v.s. others as a binary classification problem. In other words, class $k$ can be seen as class $1$ and classes $\{1,...,k-1,k+1,...,N\}$ can be seen as class $2$. The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems. Moreover, let $\widehat{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores. Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\widehat{p}(x_{i})]_{k^{\prime}}-[\widehat{p}(x_{i})]_{k}$ to sort the examples. When multiple examples have the same margin score, we further break ties by their corresponding confidence scores $[\widehat{p}(x_{i})]_{k}$. We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
@@
  Input: Pool $X$, number of rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
@@
  Initialize: Uniformly sample $B$ elements without replacement from $X$ to form $L_{0}$. Let $U_{0}\leftarrow X\backslash L_{0}$.
@@
     Initialize labeled set $L_{t}\leftarrow L_{t-1}$.
     Initialize budget per class $b\leftarrow B_{\text{train}}/2K$.
@@
        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ and break ties so that $q_{(i)}^{k}=q_{(i+1)}^{k}\implies[\widehat{p}(x_{(i)})]_{k}\geq[\widehat{p}(x_{(i+1)})]_{k}$.
        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{q^{k}_{(i)}\}_{i=1}^{N})$.
@@
     Annotate rare and uncertain examples

### v3

@@ Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods. In this work, we conduct the first study of active learning under both class imbalance and label noise. We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage classic active learning theory and methods to address issues such as batch labeling and tolerance towards label noise. We present extensive experiments on imbalanced datasets with and without label noise. Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
@@ 1 Introduction
Large-scale deep learning models are playing increasingly important roles across many industries. Human feedback and annotations have played a significant role in developing such systems. Progressively over time, we believe the role of humans in a machine learning pipeline will shift to annotating rare yet important cases. However, under data imbalance, the typical strategy of randomly choosing examples for annotation becomes especially inefficient. This is because the majority of the labeling budget would be spent on common and well-learned classes, resulting in insufficient rare class examples for training an effective model. To mitigate this issue, many recent active learning algorithms have focused on labeling more class-balanced and informative examples (Aggarwal et al., 2020; Kothawade et al., 2021; Zhang et al., 2022, 2024b; Soltani et al., 2024). For many large-scale annotation jobs, this challenge of data imbalance is further compounded by label noise – a critical and common issue that results from annotator decision fatigue and perception differences. A rich body of literature on agnostic active learning (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021) addresses this challenge on low-complexity model classes (e.g. linear models). However, for deep learning models, these algorithms often becomes ineffective due to the large model class complexity. In this paper, we propose a novel active learning strategy for both class imbalance and label noise. Our algorithm DIRECT sequentially and adaptively chooses informative and more class-balanced examples for annotation while being robust to noisy annotations. To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
@@
To bridge the gap between the imbalanced deep active learning and the agnostic active learning literature, we propose a novel reduction of the imbalanced classification problem into a set of one-dimensional agnostic active learning problems. For each class, our reduction sorts unlabeled examples into an list ordered by one-vs-rest margin scores. The objective of DIRECT is to find the optimal separation threshold which best separates the examples in the given class from the rest. By relating our problem to that of finding the best threshold classifier, we are able to employ ideas from the agnostic active learning literature to learn the separation threshold robustly under label noise. By annotating around the threshold, the annotated examples are more class-balanced and informative.
Comparing to existing active learning algorithms such as BADGE (Ash et al., 2019), Cluster-Margin (Citovsky et al., 2021), SIMILAR (Kothawade et al., 2021), GALAXY (Zhang et al., 2022) and many others, DIRECT improves significantly in label efficiencies – less annotations needed to reach the same accuracy. Notably, most existing methods mentioned above are proposed to handle batch labeling, while previous work by Zhang et al. (2022) proposes a superior performance algorithm at the cost of only allowing one annotation at a time. Our algorithm DIRECT is able to obtain the best of both worlds – practical scalability to large annotation jobs by batch labeling while also getting superior performance than all algorithms including GALAXY. On imbalanced datasets, DIRECT achieves state-of-art label efficiency on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench (Zhang et al., 2024a) framework.
@@
Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings. The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy. Both (a) and (b) are using supervised training of ResNet-18. In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2024a). $B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators. $B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY. Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world scenarios.
@@
  • We propose a novel reduction that bridges the advancement in the theoretical agnostic active learning literature to imbalanced active classification for deep neural networks.
@@
  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32). DIRECT consistently outperforms existing baseline algorithms by saving more than $60\%$ annotation cost compared to the best existing algorithm, and more than $80\%$ annotation cost compared to random sampling.
@@ 2 Related Work
Class-Balanced Deep Active Learning Active earning strategies sequentially and adaptively choose examples for annotation. Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong and Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have been shown to perform among the top when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2024a). More sophisticated methods have been proposed to optimize chosen examples’ uncertainty (Gal et al., 2017; Ducoffe and Precioso, 2018; Beluch et al., 2018), diversity (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as label noises (Khosla et al., 2022) or class imbalance Kothawade et al. (2021); Zhang et al. (2022, 2024a).
@@
Deep Active Learning under Imbalance Data imbalance and rare instances are prevalent in almost all modern machine learning applications. Active learning techniques are effective in addressing the problem in its root by collecting a more class-balanced dataset (Aggarwal et al., 2020; Kothawade et al., 2021; Emam et al., 2021; Zhang et al., 2022; Coleman et al., 2022; Jin et al., 2022; Cai, 2022; Zhang et al., 2024b). To this end, Kothawade et al. (2021) propose a submodular-based method that actively annotates examples similar to known examples of rare instances. GALAXY(Zhang et al., 2022) constructs one-dimensional linear graphs and applies graph-based active learning techniques in annotating a set of examples that are both class-balanced and uncertain. While GALAXY outperforms existing algorithms, due to a bisection procedure involved, it does not allow parallel annotation. In addition, bisection procedures are generally not robust against label noises, a prevalent challenge in real world annotation tasks. Our algorithm DIRECT mitigates all of the above shortcomings of GALAXY while outperforming it even with synchronous labeling and no label noise, beating GALAXY in its own game. Lastly, we distinguish our work from Zhang et al. (2024b), where the paper studies the algorithm selection problem. Unlike our goal of proposing a new deep active learning algorithm, the paper proposes meta algorithms to choose the right active learning algorithm among a large number of candidate algorithms.
@@
Agnostic Active Learning for Label Noise Label noise for active learning has been primarily studied under the extensive literature on agnostic learning. We refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion. All of these works, beginning with the seminal works by Balcan et al. (2006); Dasgupta et al. (2007), follow a familiar paradigm of disagreement based learning. This involves maintaining a version space of promising hypotheses at each time and constructing a disagreement region of unlabeled examples. For any unlabeled example in the disagreement region, there exists two hypotheses in the version disagreeing on their predictions. An example then chosen for annotation by sampling from a informative sampling distribution computed over the disagreement region. Several approaches have been proposed for computing such sampling distributions, e.g. Jain and Jamieson (2019); Katz-Samuels et al. (2020, 2021); Huang et al. (2015). As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al. (2021). We remark that these algorithms tend to be overly pessimistic in training deep neural nets, and this paper hopes to close this gap.
Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature. Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021). In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon. This is fundamentally different from our work, where our annotator may provide noisy labels. Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang and Chaudhuri, 2015; Chen et al., 2022). The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on such instances. In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today. Recently, Khosla et al. (2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise. Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
To our knowledge, no deep active learning literature has studied the scenario where both imbalance and label noise present. Yet, this setting is the most prevalent in real-world annotation applications.
@@ 3.1 Notations
We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation. Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown. Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted of $K$ classes. Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ of each class. We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k^{\prime}\in[K]}N_{k^{\prime}}}$.
@@
3.2 Limitations of Existing Imbalanced Active Learning Algorithms
@@
Figure 2: Demonstration of existing imbalance active learning algorithms. Ordered lists of examples are ranked by the predictive sigmoid score $\widehat{p}$. The ground truth label of each example is represented by its border – solid blue for class $1$ and dotted red for class $2$. Annotated examples are shaded.
Below we document the several active learning algorithms and how their progressive improvement. At the end, we highlight the shortcomings of the state-of-art algorithm GALAXY (Zhang et al., 2022) and motivate DIRECT’s objective of adaptively finding the optimal separation threshold. We first consider an imbalanced binary classification case, where $N_{1}<N_{2}$ without loss of generality.
Random Sampling. After annotating a significant number of examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$. This approach suffers from annotating examples that are neither class-balanced nor informative.
[…]
  1. 1.
    During active learning, the neural network is still under training and cannot perfectly separate the two classes of examples yet. Therefore, the sorted graph could have a significant number of cuts. As an example in Figure 2(b), when annotating around all of such cuts, the algorithm could waste a significant portion of the annotation budget around misclassified outliers, leading to a large number of majority class annotations.
  2. 2.
    Under label noise, the incorrect annotation could lead to more cuts in the sorted graph, further exacerbating the above issue.
  3. 3.
    GALAXY finds all cuts through a modified bisection procedure, which only allows for sequential labeling and prevents multiple annotators labeling in parallel.
In this paper, we take a DIRECT approach by identifying only the optimal separation threshold and address all of the shortcomings above.
4 A Robust Algorithm for Active Learning under Imbalance and Label Noise
@@ 4.1 An 1-D Reduction to Agnostic Active Learning
We start by considering the imbalanced binary classification setting mentioned in Section 3.2. When given a neural network model, we let $\widehat{p}:X\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores. Here, a higher sigmoid score represents a higher confidence of the example being in class $2$. We sort examples by their sigmoid predictive score similar to Section 3.2. Formally, we now define the optimal separation threshold as described in Section 3.2.
@@
Let $0=q_{(0)}\leq q_{(1)}\leq\cdots\leq q_{(N)}$, where $\{q_{(i)}\in\mathbb{R}\}_{i=1}^{N}$ is a sorted permutation of $\{\widehat{p}(x_{i})\}_{i=1}^{N}$. Further we let $\{x_{(i)}\}_{i=1}^{N}$ and $\{y_{(i)}\}_{i=1}^{N}$ denote the sorted list’s corresponding examples and labels. We define the optimal separation threshold as $j^{\star}\in\{0,1,...,N\}$ such that
@@
$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=1:i>j\}\rvert\right).$ (1)
@@
In other words, on either side of $j^{\ast}$, it has the largest discrepancy in the number of examples between the two classes. This captures the intuition of Figure 2(a) — our goal is to find a threshold that best separates one class from the other. We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
@@
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}$ is defined as $h_{j}(q)=\begin{cases}1&\text{if}\,q\leq q_{(j)}\\ 2&\text{if}\,q>q_{(j)}\end{cases}$. Here, $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times. The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1). Namely, with ties broken similar to above, $j^{\star}=\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$.
Figure 3: Visualization of multi-class classification. For each class, we formulate the problem as a one-vs-rest binary classification problem by sorting examples based on margin scores. The black arrows indicates the optimal separation thresholds for each class.
@@
Multi-Class Classification. To generalize the above problem formulation to multi-class classification, we follow a similar strategy to Zhang et al. (2022). As shown in Figure 3, for each class $k$, we can view the problem of class-$k$ v.s. others as a binary classification problem. The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems. Moreover, let $\widetilde{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function, mapping examples to softmax scores. For each class $k$, we use the margin scores $\widehat{p}_{i}^{k}:=[\widetilde{p}(x_{i})]_{k}-\max_{k^{\prime}}[\widetilde{p}(x_{i})]_{k^{\prime}}$ to sort the examples and break ties by their corresponding confidence scores $[\widetilde{p}(x_{i})]_{k}$. Formally,
$\left(q_{(1)}^{k}\leq\mkern-2.0mu{\cdots}\leq q_{(N)}^{k}\mkern-1.0mu{:}\text{ sorted permutation of\,}\{\widehat{p}_{i}^{k}\}_{i=1}^{N}\right)\mkern-1.0mu{\land}\left(q_{(i)}^{k}\mkern-1.0mu{=}q_{(i+1)}^{k}\Rightarrow[\widetilde{p}(x_{i})]_{k}\geq[\widetilde{p}(x_{i+1})]_{k}\right).$ (2)
Note that sorting by margin scores is equivalent to sorting by sigmoid scores for binary classification.
@@
  Input: Pool $X$, #Rounds $T$, retraining batch size $B_{\text{train}}$, number of parallel annotations $B_{\text{parallel}}$.
@@
  Initialize: Uniformly sample $B$ elements from $X$ to form $L_{0}$. Let $U_{0}\leftarrow X\backslash L_{0}$.
@@
     Initialize labeled set $L_{t}\leftarrow L_{t-1}$ and budget per class $b\leftarrow B_{\text{train}}/2K$.
@@
        Sort margin scores $0=q_{(0)}^{k}\leq q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ based on equation (2).
        Let $x_{(i)}^{k},y_{(i)}^{k}$ denote the example and label corresponding to $q_{(i)}^{k}$.
        Identify threshold for class $k$: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x^{k}_{(i)},y^{k}_{(i)}),q^{k}_{(i)}\}_{i=1}^{N})$.
@@
     Annotate examples around the identified threshold